Log(10+9x)-log(x-1)=2
Answer by jim_thompson5910(35256)

I'll get you started.

log(10+9x)-log(x-1)=2

log[(10+9x)/(x-1)]=2

(10+9x)/(x-1) = 10^2

(10+9x)/(x-1) = 100

10+9x = 100(x-1)

I'll let you finish up. Don't forget to check all possible answers you get.